OpenStudy (anonymous):

Sean’s Bikes has the same number of bicycles as tricycles in stock, but they have 54 more wheels than seats. What is the total number of bicycles and tricycles in the shop? Please explain

OpenStudy (mathmagician):

Let x be tricycles and y b bicycles. You know, that x=y And 54+x+y=3x+2y (3x+2y is number of wheels and 54+x+y the number of seats in the shop). Since x=y, you have 54+2x=3x+2x and x=18.

OpenStudy (anonymous):

Thanks!